# presep= ]e1,e2,e3

### v1

@@ Appendix B Proofs of results from Section 4
$\sqrt{k_{n}h_{n}}\left|\frac{\overline{F}^{x}(tu_{n})}{\overline{F}^{x}(u_{n})}-t^{-\alpha_{x}}\right|\leq\text{cst}\ t^{-\alpha_{x}}\left(1+\log(t)\right)\sqrt{k_{n}h_{n}}\frac{n}{k_{n}}\frac{u_{n}^{-\alpha_{x}}\log(u_{n})}{\log(u_{n})}$
$=\text{cst}\ t^{-\alpha_{x}}\left(1+\log(t)\right)\sqrt{k_{n}h_{n}}\frac{n}{k_{n}}\frac{\frac{k_{n}}{n}+\mathcal{O}\left(u_{n}^{-\beta_{x}}\right)}{\log(u_{n})}$
$=\text{cst}\ t^{-\alpha_{x}}\left(1+\log(t)\right)\left(\frac{\sqrt{k_{n}h_{n}}}{\log(u_{n})}+\frac{\mathcal{O}\left(n\sqrt{h_{n}}u_{n}^{-\beta_{x}}\right)}{\log(u_{n})\sqrt{k_{n}}}\right).$
The mapping $t\mapsto t^{-\alpha_{x}-1}\left(1+\log(t)\right)$ is integrable on $(1,\infty)$ and hence it remains to examine the deterministic sequence
$\left(\frac{\sqrt{k_{n}h_{n}}}{\log(u_{n})}+\frac{\mathcal{O}\left(n\sqrt{h_{n}}u_{n}^{-\beta_{x}}\right)}{\log(u_{n})\sqrt{k_{n}}}\right)=\frac{\sqrt{k_{n}h_{n}}}{\log(u_{n})}\left(1+\frac{n}{k_{n}u_{n}^{\beta_{x}}}\right).$
Notice first that
$\frac{n}{k_{n}u_{n}^{\beta_{x}}}=\frac{1}{\overline{F}^{x}(u_{n})u_{n}^{\beta_{x}}}\to 0,$
as $\beta_{x}>\alpha_{x}$.
As argued in the proof of Theorem 7 we have $\log\left(n/k_{n}\right)\sim\alpha_{x}\log(u_{n})$ and thus
$\frac{\sqrt{k_{n}h_{n}}}{\log(u_{n})}\left(1+\frac{n}{k_{n}u_{n}^{\beta_{x}}}\right)\sim\frac{\alpha_{x}\sqrt{k_{n}h_{n}}}{\log(n/k_{n})}\left(1+\frac{n}{k_{n}u_{n}^{\beta_{x}}}\right)\to 0.$
@@
$\sqrt{k_{n}h_{n}}B_{n}^{x}(s_{0})$ $\leq\mathcal{O}(\sqrt{k_{n}h_{n}}h_{n})+\mathcal{O}\left(\sqrt{k_{n}h_{n}}\log\left(u_{n}\right)h_{n}\right)+\mathcal{O}\left(\frac{\sqrt{k_{n}h_{n}}}{\log\left(\frac{n}{k_{n}}\right)}\right)$
$=\mathcal{O}\left(\sqrt{k_{n}h_{n}^{3}}\log\left(\frac{n}{k_{n}}\right)\right)+\mathcal{O}\left(\frac{\sqrt{k_{n}h_{n}}}{\log\left(\frac{n}{k_{n}}\right)}\right)\to 0.$
@@
$\sum_{j=m}^{r_{n}}\frac{\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}\right\}}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>tu_{n}\right\}}\right]}{h_{n}\overline{F}^{x}(u_{n})}$
$\quad=\sum_{j=m}^{r_{n}}\frac{\text{cov}\left(K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}\right\}}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>tu_{n}\right\}}\right)}{h_{n}\overline{F}^{x}(u_{n})}$
$\qquad+(r_{n}-m)h_{n}\overline{F}^{x}(u_{n})\frac{\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}\right\}}\right]}{h_{n}\overline{F}^{x}(u_{n})}\frac{\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>tu_{n}\right\}}\right]}{h_{n}\overline{F}^{x}(u_{n})},$
@@
$\sum_{j=m}^{r_{n}}\frac{\text{cov}\left(K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}\right\}}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>tu_{n}\right\}}\right)}{h_{n}\overline{F}^{x}(u_{n})}\leq\frac{\text{cst}}{h_{n}\overline{F}^{x}(u_{n})}\left(h_{n}^{2-c_{x}}\overline{F}^{x}(u_{n})+h_{n}^{c_{x}(\eta-1)}\right)$
$\quad=\text{cst}(h_{n}^{1-c_{x}}+h_{n}^{c_{x}(\eta-1)-1}\overline{F}^{x}(u_{n})^{-1})\to 0,$
as $n\to\infty$, for some constant $c_{x}\geq 0$, by assumption. This yields condition 2.
We proceed to establish the logarithmic anti-clustering condition 6, focusing, as above, on the covariance rather than the cross-term. Let $q>2$ and $p>1$ such that $2/q=1-1/p$. Rio’s covariance inequality and calculations similar to those in the proof of Theorem 5 yield
$\left|\text{cov}\left(K\left(\frac{x-X_{0}}{h_{n}}\right)\log_{+}\left(\frac{Y_{0}}{u_{n}A}\right),K\left(\frac{x-X_{j}}{h_{n}}\right)\log_{+}\left(\frac{Y_{j}}{u_{n}A}\right)\right)\right|$
$\leq 2p(2\alpha(j))^{1/p}\mathbb{E}\left[K^{q}\left(\frac{x-X_{0}}{h_{n}}\right)\log_{+}^{q}\left(\frac{Y_{0}}{u_{n}A}\right)\right]^{2/q}=\mathcal{O}(\alpha(j)^{1/p}h_{n}^{2/q}\overline{F}^{x}(u_{n})^{2/q}).$
@@
$\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)K\left(\frac{x-X_{j}}{h_{n}}\right)\log_{+}\left(\frac{Y_{0}}{u_{n}A}\right)\log_{+}\left(\frac{Y_{j}}{u_{n}A}\right)\right]$
$=h_{n}^{2}\overline{F}^{x}(u_{n})\int_{[-1,1]^{2}}\biggl{\{}\frac{\mathbb{E}\left[\log_{+}\left(\frac{Y_{0}}{u_{n}A}\right)\log_{+}\left(\frac{Y_{j}}{u_{n}A}\right)\mid X_{0}=x+h_{n}w_{1},X_{j}=x+h_{n}w_{2}\right]}{\overline{F}^{x}(u_{n})}$
×$\qquad\qquad\qquad\qquad\qquad\qquad\times g(x+w_{1}h_{n},x+w_{2}h_{n})K\left(w_{1}\right)K\left(w_{2}\right)\biggr{\}}\mathrm{d}(w_{1},w_{2})=\mathcal{O}\left(h_{n}^{2}\overline{F}^{x}(u_{n})\right),$
@@
$\frac{1}{h_{n}\overline{F}^{x}(u_{n})}\left|\sum_{j=m}^{r_{n}}\text{cov}\left(K\left(\frac{x-X_{0}}{h_{n}}\right)\log_{+}\left(\frac{Y_{0}}{u_{n}A}\right),K\left(\frac{x-X_{j}}{h_{n}}\right)\log_{+}\left(\frac{Y_{j}}{u_{n}A}\right)\right)\right|$
$\leq\frac{\text{cst}}{h_{n}\overline{F}^{x}(u_{n})}\left(\sum_{j=1}^{d_{n}-1}h_{n}^{2}\overline{F}^{x}(u_{n})+\sum_{j=d_{n}}^{\infty}h_{n}^{2/q}\overline{F}^{x}(u_{n})^{2/q}j^{-\eta/p}\right)$
$\leq\text{cst}\left(d_{n}h_{n}+d_{n}^{-\eta/p+1}h_{n}^{2/q-1}\overline{F}^{x}(u_{n})^{2/q-1}\right)\sim\text{cst}\left(h_{n}^{1-c^{\prime}_{x}}+h_{n}^{-1/p+c^{\prime}_{x}(\eta/p-1)}\overline{F}^{x}(u_{n})^{-1/p}\right)\to 0,$
where as above, we have $d_{n}=\lceil h_{n}^{-c^{\prime}_{x}}\rceil$ for some $c^{\prime}_{x}\geq 0$. ∎
@@
Theorem 6 ensures that the first two limit conditions for consistency are fulfilled for $h_{n}\asymp\overline{F}^{x}(u_{n})^{e}$ for $e>0$ and $k_{n}\asymp n^{\delta}$ for $\delta\in(e/(1+e),1)$. Let $r_{n}\asymp n^{a},\ a\in(0,1)$. Then $r_{n}/n\to 0$ and
$r_{n}h_{n}\overline{F}^{x}(u_{n})\asymp n^{a}\left(\frac{k_{n}}{n}\right)^{1+e}\asymp n^{a-(1-\delta)(1+e)}\to 0,$
since $(1-\delta)(1+e)>1$. Let $\ell_{n}\asymp n^{\psi},\ \psi\in(0,a)$. Then $1/\ell_{n}\to 0$, $\ell_{n}/r_{n}\to 0$ and
@@
since for every $\nu>0$ there exists $a\in(0,1)$ and $\psi\in(0,a)$ such that $01-a-\nu\psi<$ 0. This concludes the second limit condition. Lemma 16 yields that the third condition holds if $\lim_{n\to\infty}\sqrt{\overline{F}^{x}(u_{n})nh_{n}}h_{n}^{2}=0$. This is indeed true since
$\sqrt{\overline{F}^{x}(u_{n})nh_{n}}h_{n}^{2}\asymp\sqrt{n\left(\frac{k_{n}}{n}\right)^{1+5e}}\asymp\sqrt{n^{1-(1+5e)(1-\delta)}},$
which vanishes asymptotically for $\delta<e/(e+1/3)<e/(e+1/5)$. As we already require $\delta>e/(1+e)$, we additionally require $1/4<e$. According to Lemma 1 it suffices to check the limits (8) in order ensure the last two of the above five conditions. Leet $\delta\in(e/(1+e),e/(e+1/3))$. Then
$\sqrt{k_{n}h_{n}^{3}}\log\left(\frac{n}{k_{n}}\right)\asymp\sqrt{n^{\delta-3e(1-\delta)}}\log(n^{1-\delta})\to 0.$
@@
$\delta\in\left(\frac{e}{1+e},\frac{e}{\frac{1}{3}+e}\wedge\frac{\frac{e}{2}+\frac{\beta_{x}}{\alpha_{x}}-1}{\frac{e}{2}+\frac{\beta_{x}}{\alpha_{x}}-\frac{1}{2}}\right),$
and note that this is a valid interval for $1/4<e$, since $\beta_{x}/\alpha_{x}>1$, by assumption. It then holds that
$\sqrt{h_{n}}k_{n}^{\frac{\beta_{x}}{\alpha_{x}}-\frac{1}{2}}n^{1-\frac{\beta_{x}}{\alpha_{x}}}\asymp\sqrt{n^{-e(1-\delta)}}n^{\delta(\frac{\beta_{x}}{\alpha_{x}}-\frac{1}{2})+1-\frac{\beta_{x}}{\alpha_{x}}}=n^{\delta(\frac{e}{2}+\frac{\beta_{x}}{\alpha_{x}}-\frac{1}{2})-\left(\frac{e}{2}+\frac{\beta_{x}}{\alpha_{x}}-1\right)}\to 0,$
@@
According to Lemma 3, the anti-clustering conditions follow if (12) holds. The first limit expression of (12) is given by
$h_{n}^{1-c_{x}}+h_{n}^{c_{x}(\eta-1)-1}\overline{F}^{x}(u_{n})^{-1}\asymp\overline{F}^{x}(u_{n})^{e(1-c_{x})}+\overline{F}^{x}(u_{n})^{ec_{x}(\eta-1)+(e+1)}.$
We require that $e>(\eta-2)^{-1}$. Since $\eta>2$, there exists $c_{x}\in(0,1)$, such that the exponents above are positive, and consequently that the expression converges to zero as $n\to\infty$. For the second limit expression of (12), note that
$h_{n}^{1-c^{\prime}_{x}}+h_{n}^{-1/p+c^{\prime}_{x}(\eta/p-1)}\overline{F}^{x}(u_{n})^{-1/p}\asymp\overline{F}^{x}(u_{n})^{e(1-c^{\prime}_{x})}+\overline{F}^{x}(u_{n})^{ec^{\prime}_{x}(\eta/p-1)+(e+1)/p}.$
Since $p>1$ and $e>(\eta-2)^{-1}$, we can find $c^{\prime}_{x}\in(0,1)$ such that $\eta-(c^{\prime}_{x})^{-1}-(ec^{\prime}_{x})^{-1}>1$, and consequently a $p$ such that $1<p<\eta-(c^{\prime}_{x})^{-1}-(ec^{\prime}_{x})^{-1}$. For such constants the exponents above are positive, and consequently that the expression converges to zero as $n\to\infty$. ∎
@@
It holds that $k_{n}/n\asymp\log(n)^{2}/n\to 0$ and hence $\lim_{n\to\infty}u_{n}=\infty$. Secondly
$nh_{n}\overline{F}^{x}(u_{n})=k_{n}h_{n}\asymp\log(n)^{2-e}\to\infty.$
As argued in the proof of Theorem 7, we have $\log\left(n/k_{n}\right)\sim\alpha_{x}\log(u_{n})$ and thus
$h_{n}\log(u_{n})\sim\frac{h_{n}\log(n/k_{n})}{\alpha_{x}}\leq\text{cst}\log(n)^{1-e}\to 0.$
@@
by arguments similar to the proof of Theorem 8, which concludes the second normality limit condition above. Note also that
$\sqrt{\overline{F}^{x}(u_{n})nh_{n}}h_{n}^{2}\asymp\sqrt{k_{n}h_{n}}h_{n}^{2}\asymp\log(n)^{1-5e/2}\to 0,$
@@
$\frac{\sqrt{k_{n}h_{n}}}{\log\left(\frac{n}{k_{n}}\right)}\sim\frac{\sqrt{k_{n}h_{n}}}{\log\left(n\right)}\asymp\frac{\sqrt{\log\left(n\right)^{2-e}}}{\log\left(n\right)}=\log\left(n\right)^{-e/2}\to 0,$
@@
$\sqrt{k_{n}h_{n}}h_{n}\log(u_{n})\sim\frac{\sqrt{k_{n}h_{n}}h_{n}\log(n/k_{n})}{\alpha_{x}}\leq\text{cst}\sqrt{k_{n}h_{n}}h_{n}\log(n)\asymp\log\left(n\right)^{2-3e/2}\to 0.$

### v2

@@ Appendix B Proofs of results from Section 4
$\sqrt{k_{n}h_{n}}\left|\frac{\overline{F}^{x}(tu_{n}^{x})}{\overline{F}^{x}(u_{n}^{x})}-t^{-1/\gamma(x)}\right|\leq\text{cst}\ t^{-1/\gamma(x)}\left(1+\log(t)\right)\sqrt{k_{n}h_{n}}\frac{n}{k_{n}}\frac{(u_{n}^{x})^{-1/\gamma(x)}\log(u_{n}^{x})}{\log(u_{n}^{x})}$
$=\text{cst}\ t^{-1/\gamma(x)}\left(1+\log(t)\right)\sqrt{k_{n}h_{n}}\frac{n}{k_{n}}\frac{\frac{k_{n}}{n}+\mathcal{O}\left((u_{n}^{x})^{-\zeta_{x}}\right)}{\log(u_{n}^{x})}$
$=\text{cst}\ t^{-1/\gamma(x)}\left(1+\log(t)\right)\left(\frac{\sqrt{k_{n}h_{n}}}{\log(u_{n}^{x})}+\frac{\mathcal{O}\left(n\sqrt{h_{n}}(u_{n}^{x})^{-\zeta_{x}}\right)}{\log(u_{n}^{x})\sqrt{k_{n}}}\right).$
The mapping $t\mapsto t^{-1/\gamma(x)-1}\left(1+\log(t)\right)$ is integrable on $(1,\infty)$ and hence it remains to examine the deterministic sequence
$\left(\frac{\sqrt{k_{n}h_{n}}}{\log(u_{n}^{x})}+\frac{\mathcal{O}\left(n\sqrt{h_{n}}(u_{n}^{x})^{-\zeta_{x}}\right)}{\log(u_{n}^{x})\sqrt{k_{n}}}\right)=\frac{\sqrt{k_{n}h_{n}}}{\log(u_{n}^{x})}\left(1+\frac{n}{k_{n}(u_{n}^{x})^{\zeta_{x}}}\right).$
Notice first that, since $\zeta_{x}\gamma(x)>1$,
$\frac{n}{k_{n}(u_{n}^{x})^{\zeta_{x}}}=\frac{1}{\overline{F}^{x}(u_{n}^{x})(u_{n}^{x})^{\zeta_{x}}}\to 0,$
As argued in the proof of Theorem 7 we have $\log\left(n/k_{n}\right)\sim(1/\gamma(x))\log(u_{n}^{x})$ and thus
$\frac{\sqrt{k_{n}h_{n}}}{\log(u_{n}^{x})}\left(1+\frac{n}{k_{n}(u_{n}^{x})^{\zeta_{x}}}\right)\sim\frac{\sqrt{k_{n}h_{n}}}{\gamma(x)\log(n/k_{n})}\left(1+\frac{n}{k_{n}(u_{n}^{x})^{\zeta_{x}}}\right)\to 0.$
@@
$\sqrt{k_{n}h_{n}}B_{n}^{x}(s_{0})$ $\leq\mathcal{O}(\sqrt{k_{n}h_{n}}h_{n})+\mathcal{O}\left(\sqrt{k_{n}h_{n}}\log\left(u_{n}^{x}\right)h_{n}\right)+\mathcal{O}\left(\frac{\sqrt{k_{n}h_{n}}}{\log\left(n/k_{n}\right)}\right)$
$=\mathcal{O}\left(\sqrt{k_{n}h_{n}^{3}}\log\left(n/k_{n}\right)\right)+\mathcal{O}\left(\frac{\sqrt{k_{n}h_{n}}}{\log\left(n/k_{n}\right)}\right)\to 0.$
@@
$\sum_{j=m}^{r_{n}}\frac{\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}^{x}\right\}}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>tu_{n}^{x}\right\}}\right]}{h_{n}\overline{F}^{x}(u_{n}^{x})}$
$\quad=\sum_{j=m}^{r_{n}}\frac{\text{cov}\left(K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}^{x}\right\}}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>tu_{n}^{x}\right\}}\right)}{h_{n}\overline{F}^{x}(u_{n}^{x})}$
$\qquad+(r_{n}-m)h_{n}\overline{F}^{x}(u_{n}^{x})\frac{\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}^{x}\right\}}\right]}{h_{n}\overline{F}^{x}(u_{n}^{x})}\frac{\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>tu_{n}^{x}\right\}}\right]}{h_{n}\overline{F}^{x}(u_{n}^{x})},$
@@
$\sum_{j=m}^{r_{n}}\frac{\text{cov}\left(K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}^{x}\right\}}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>tu_{n}^{x}\right\}}\right)}{h_{n}\overline{F}^{x}(u_{n}^{x})}\leq\frac{\text{cst}}{h_{n}\overline{F}^{x}(u_{n}^{x})}\left(h_{n}^{2-c_{x}}\overline{F}^{x}(u_{n}^{x})+h_{n}^{c_{x}(\eta-1)}\right)$
$\quad=\text{cst}(h_{n}^{1-c_{x}}+h_{n}^{c_{x}(\eta-1)-1}\overline{F}^{x}(u_{n}^{x})^{-1})\to 0,$
as $n\to\infty$, for some constant $c_{x}>0$, by assumption. This yields condition 2.
We proceed to establish the logarithmic anti-clustering condition 6, focusing, as above, on the covariance rather than the cross-term. Let $q>2$ and $p>1$ such that 2/q=11/p2/q=1-1/p. Rio’s covariance inequality and calculations similar to those in the proof of Theorem 5 yield
$\left|\text{cov}\left(K\left(\frac{x-X_{0}}{h_{n}}\right)\log_{+}\left(\frac{Y_{0}}{u_{n}^{x}A}\right),K\left(\frac{x-X_{j}}{h_{n}}\right)\log_{+}\left(\frac{Y_{j}}{u_{n}^{x}A}\right)\right)\right|$
$\leq 2p(2\alpha(j))^{1/p}\mathbb{E}\left[K^{q}\left(\frac{x-X_{0}}{h_{n}}\right)\log_{+}^{q}\left(\frac{Y_{0}}{u_{n}^{x}A}\right)\right]^{2/q}=\mathcal{O}(\alpha(j)^{1/p}h_{n}^{2/q}\overline{F}^{x}(u_{n}^{x})^{2/q}).$
@@
$\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)K\left(\frac{x-X_{j}}{h_{n}}\right)\log_{+}\left(\frac{Y_{0}}{u_{n}^{x}A}\right)\log_{+}\left(\frac{Y_{j}}{u_{n}^{x}A}\right)\right]$
$=h_{n}^{2}\overline{F}^{x}(u_{n}^{x})\int_{[-1,1]^{2}}\biggl\{\frac{\mathbb{E}\left[\log_{+}\left(\frac{Y_{0}}{u_{n}^{x}A}\right)\log_{+}\left(\frac{Y_{j}}{u_{n}^{x}A}\right)\mid X_{0}=x+h_{n}w_{1},X_{j}=x+h_{n}w_{2}\right]}{\overline{F}^{x}(u_{n}^{x})}$
$\qquad\qquad\qquad\qquad\qquad\qquad\times g(x+w_{1}h_{n},x+w_{2}h_{n})K\left(w_{1}\right)K\left(w_{2}\right)\biggr\}\mathrm{d}(w_{1},w_{2})=\mathcal{O}\left(h_{n}^{2}\overline{F}^{x}(u_{n}^{x})\right),$
@@
$\frac{1}{h_{n}\overline{F}^{x}(u_{n}^{x})}\left|\sum_{j=m}^{r_{n}}\text{cov}\left(K\left(\frac{x-X_{0}}{h_{n}}\right)\log_{+}\left(\frac{Y_{0}}{u_{n}^{x}A}\right),K\left(\frac{x-X_{j}}{h_{n}}\right)\log_{+}\left(\frac{Y_{j}}{u_{n}^{x}A}\right)\right)\right|$
$\leq\frac{\text{cst}}{h_{n}\overline{F}^{x}(u_{n}^{x})}\left(\sum_{j=1}^{d_{n}-1}h_{n}^{2}\overline{F}^{x}(u_{n}^{x})+\sum_{j=d_{n}}^{\infty}h_{n}^{2/q}\overline{F}^{x}(u_{n}^{x})^{2/q}j^{-\eta/p}\right)$
$\leq\text{cst}\left(d_{n}h_{n}+d_{n}^{-\eta/p+1}h_{n}^{2/q-1}\overline{F}^{x}(u_{n}^{x})^{2/q-1}\right)\sim\text{cst}\left(h_{n}^{1-c^{\prime}_{x}}+h_{n}^{-1/p+c^{\prime}_{x}(\eta/p-1)}\overline{F}^{x}(u_{n}^{x})^{-1/p}\right)\to 0,$
where as above, we have $d_{n}=\lceil h_{n}^{-c^{\prime}_{x}}\rceil$ for some $c^{\prime}_{x}>0$. ∎
@@
Theorem 6 ensures that the first two limit conditions for consistency are fulfilled for $k_{n}\asymp n^{\delta}$ for $\delta\in(0,3\nu/(2(1+\nu)))$ and $h_{n}\asymp\overline{F}^{x}(u_{n}^{x})^{e}$ for $e\in(\delta/(3(1-\delta)),\delta/(1-\delta))$. Let $r_{n}\asymp n^{a},\ a=1-2\delta/3$. Then $r_{n}/n\to 0$ and
$r_{n}h_{n}\overline{F}^{x}(u_{n}^{x})\asymp n^{a}\left(\frac{k_{n}}{n}\right)^{1+e}\asymp n^{a-(1-\delta)(1+e)}\to 0,$
since $(1-\delta)(1+e)>1-2\delta/3$. Let $\ell_{n}\asymp n^{\psi},\ \psi\in(0,a)$. Then $1/\ell_{n}\to 0$, $\ell_{n}/r_{n}\to 0$ and
@@
since upper bound on $\delta$ ensures that for every $\nu>0$ there exists a small $\varepsilon>0$ such that for $\psi=a-\varepsilon$, it holds that $1-a-\nu\psi<0$. This concludes the second limit condition. Lemma 16 yields that the third condition holds if $\lim_{n\to\infty}\sqrt{\overline{F}^{x}(u_{n}^{x})nh_{n}}h_{n}^{2}=0$. This is indeed true since
$\sqrt{\overline{F}^{x}(u_{n}^{x})nh_{n}}h_{n}^{2}\asymp\sqrt{n\left(\frac{k_{n}}{n}\right)^{1+5e}}\asymp\sqrt{n^{1-(1+5e)(1-\delta)}},$
which vanishes asymptotically for $\delta/(5(1-\delta))<\delta/(3(1-\delta))<e$. According to Lemma 1 it suffices to check the limits of equation (5) in order ensure the last two of the above five conditions. The lower bound on $e$ ensures
$\sqrt{k_{n}h_{n}^{3}}\log\left(n/k_{n}\right)\asymp\sqrt{n^{\delta-3e(1-\delta)}}\log(n^{1-\delta})\to 0.$
@@
$\delta\in\left(0,\ \frac{3}{2}\frac{\nu}{1+\nu}\wedge\frac{2(\zeta_{x}-1/\gamma(x))}{2\zeta_{x}-1/\gamma(x)}\right),$
where the latter fraction of the upper bound of $\delta$ ensures that $e$ is in a non-empty interval, since $\zeta_{x}\gamma(x)>1$, by assumption. It then holds that
$\sqrt{h_{n}}k_{n}^{\zeta_{x}\gamma(x)-\frac{1}{2}}n^{1-\zeta_{x}\gamma(x)}\asymp\sqrt{h_{n}}n^{\delta(\zeta_{x}\gamma(x)-\frac{1}{2})+1-\zeta_{x}\gamma(x)}\to 0,$
@@
According to Lemma 3, the anti-clustering conditions follow if equation (9) holds. The first limit expression of equation (9) is given by
$h_{n}^{1-c_{x}}+h_{n}^{c_{x}(\eta-1)-1}\overline{F}^{x}(u_{n}^{x})^{-1}\asymp\overline{F}^{x}(u_{n}^{x})^{e(1-c_{x})}+\overline{F}^{x}(u_{n}^{x})^{ec_{x}(\eta-1)+(e+1)}.$
Since $\eta>2$, there exists $c_{x}\in(0,1)$, such that the exponents above are positive, and consequently that the expression converges to zero as $n\to\infty$. For the second limit expression of equation (9), note that
$h_{n}^{1-c^{\prime}_{x}}+h_{n}^{-1/p+c^{\prime}_{x}(\eta/p-1)}\overline{F}^{x}(u_{n}^{x})^{-1/p}\asymp\overline{F}^{x}(u_{n}^{x})^{e(1-c^{\prime}_{x})}+\overline{F}^{x}(u_{n}^{x})^{ec^{\prime}_{x}(\eta/p-1)+(e+1)/p}.$
For $c^{\prime}_{x}\in(0,1)$ and $p\in(1,(1+e)/(ec^{\prime}_{x})+\eta)$ the exponents above are positive, and consequently the expression converges to zero as $n\to\infty$. ∎
@@
It holds that $k_{n}/n\asymp\log(n)^{2}/n\to 0$ and hence $\lim_{n\to\infty}u_{n}^{x}=\infty$. Secondly
$nh_{n}\overline{F}^{x}(u_{n}^{x})=k_{n}h_{n}\asymp\log(n)^{2-e}\to\infty.$
As argued in the proof of Theorem 7, we have $\log\left(n/k_{n}\right)\sim(1/\gamma(x))\log(u_{n}^{x})$ and thus
$h_{n}\log(u_{n}^{x})\sim\gamma(x)h_{n}\log(n/k_{n})\leq\text{cst}\log(n)^{1-e}\to 0.$
@@
since for every $\nu>0$ there exists $a\in(0,1)$ and $\psi\in(0,a)$ such that $1-a-\nu\psi<0$. This concludes the second normality limit condition above. Note also that
$\sqrt{\overline{F}^{x}(u_{n}^{x})nh_{n}}h_{n}^{2}\asymp\sqrt{k_{n}h_{n}}h_{n}^{2}\asymp\log(n)^{1-5e/2}\to 0,$
@@
$\frac{\sqrt{k_{n}h_{n}}}{\log\left(n/k_{n}\right)}\sim\frac{\sqrt{k_{n}h_{n}}}{\log\left(n\right)}\asymp\frac{\sqrt{\log\left(n\right)^{2-e}}}{\log\left(n\right)}=\log\left(n\right)^{-e/2}\to 0,$
@@
$\sqrt{k_{n}h_{n}}h_{n}\log(u_{n}^{x})\sim\gamma(x)\sqrt{k_{n}h_{n}}h_{n}\log(n/k_{n})\leq\text{cst}\sqrt{k_{n}h_{n}}h_{n}\log(n)\asymp\log\left(n\right)^{2-3e/2}\to 0.$